OpenStudy (anonymous):

3x-6y+2=0 Im stuck trying to find the slope and y int

OpenStudy (anonymous):

try changing this equation to slope-int form: y=mx+b , where m is the slope and b is the y intercept. so \(3x-6y+2=0\\-6y=-3x-2\\6y=3x+2\\y=\frac{3}{6}x+\frac{2}{6}\\y=\frac{1}{2}x+\frac{1}{3}\) therefore m=1/2 and y-int is (0, 1/3)

OpenStudy (anonymous):

So in the future you basically re arrange the problem to have Y and X on separate sides?

OpenStudy (anonymous):

@Data_LG2 Mind helping me with one more?

OpenStudy (anonymous):

you just have to follow the slope-int form \(y=mx+b\) :) sure!

OpenStudy (anonymous):

Ok. y-5=2/5(x-1)

OpenStudy (anonymous):

Just do the same thing like I did before. Try it, I'll check if you got it :)

OpenStudy (anonymous):

use the "draw" button so that i can see your solution

OpenStudy (anonymous):

I already worked it out on paper before and got y=2/5x-2/5

OpenStudy (anonymous):

you are almost there! but you forgot to transpose "5" on the right side

OpenStudy (anonymous):

Which would make it? Im pretty bad with math

OpenStudy (anonymous):

your orignal question is  simplify it and tell me what you get. It's okay I understand, everyone is bad in math. Just practice solving problems and you'll be good with it :)